# Cost-Sensitive Learning to Defer to Multiple Experts with Workload Constraints

### v2

@@ C.3 Expert Properties
$\begin{split}\mbox{FPR}_{j}=\frac{{\mathbb{E}[C]_{j}}}{\lambda(1-\mathbb{P}(\mbox{y}_{i}=1))}-\frac{\mathbb{P}(\mbox{y}_{i}=1)}{\lambda(1-\mathbb{P}(\mbox{y}_{i}=1))}\mbox{FNR}_{j},\end{split}$ (25)
@@ Appendix D OvA Classifiers and HEM
We model both the OvA Classifiers and the HEM with LightGBM [Ke et al., 2017] classifiers. These models are trained on the alerts raised by the Alert Model, and the corresponding expert predictions, ranging from months four to six of the BAF dataset. Both methods are trained by minimizing the weighted log-loss. The choice of hyperparameters is defined through Bayesian search [Akiba et al., 2019] on an extensive grid, for 120 total trials, with 100 start-up trials, with validation done on the seventh month. The hyperparameter search space is detailed in Table 7.

### v3

@@ C.3 Expert Properties
$\begin{split}\mbox{FPR}_{j}=\frac{{\mathbb{E}[C]_{j}}}{\lambda(1-\mathbb{P}(y_{i}=1))}-\frac{\mathbb{P}(y_{i}=1)}{\lambda(1-\mathbb{P}(y_{i}=1))}\mbox{FNR}_{j},\end{split}$ (30)
@@ Appendix D OvA Classifiers and HEM
We model both the OvA Classifiers and the HEM with LightGBM (Ke et al., 2017) classifiers. These models are trained on the alerts raised by the Alert Model, and the corresponding expert predictions, ranging from months four to six of the BAF dataset. Both methods are trained by minimizing the weighted log-loss. The choice of hyperparameters is defined through Bayesian search (Akiba et al., 2019) on an extensive grid, for 120 total trials, with 100 start-up trials, with validation done on the seventh month. The hyperparameter search space is detailed in Table 7.